# Propagation of Precessing Jet in Envelope of Tidal Disruption Events

### v1

@@ 4.3 Comparison with Observations
    AT2020ocn is an optical transient from the centre of a previously quiescent galaxy (Gezari et al., 2020). Subsequently, the Neutron Star Interior Composition Explorer (NICER) started to observe and found that, unlike the optical and ultraviolet emission, the soft X-ray radiation exhibited clear quasi-periodic modulations with a period of $\sim 15$ days (Pasham et al., 2024). In our model framework, the envelope of AT2020ocn has a large inclination angle, and the jet precession angle is smaller than the envelope inclination. As a result, the jet is choked within the envelope, forming a cocoon. The cocoon produces thermal X-ray emission during shock breakout and the subsequent cooling. Besides, similar modulations are not present in the optical bands (Pasham et al., 2024), which suggests that the optical radiation likely originates primarily from thermal emission at the envelope surface.
@@ 5 Conclusion
    The envelope density along its polar regions is significantly reduced, allowing the jets to escape freely from these funnels. For $\theta_{\rm LT}<\theta_{\rm env}$, observers with line of sight along these free zones can, therefore, detect a strong episodic jet, as the X-ray quasi-perodic dip shown in J1644. Away from these free zones, jets will be choked in most cases, and can only break out successfully at small precession angle. An off-axis jetted TDE, like J12580, is possible when looking along the choked zone.
@@
We are very grateful to Xiangyu Wang, and Rongfeng Shen for their helpful discussions. This work is supported by the National Natural Science Foundation of China under grant 12473012and the National Key R&D Program of China (Nos. 2020YFC2201400, SQ2023YFC220007). W.H.Lei. acknowledges support by the science research grants from the China Manned Space Project with NO.CMS-CSST-2021-B11.

### v2

@@ 4.3 Comparison with Observations
    AT2020ocn is an optical transient from the centre of a previously quiescent galaxy (Gezari et al., 2020). Subsequently, the Neutron Star Interior Composition Explorer (NICER) started to observe and found that, unlike the optical and ultraviolet emission, the soft X-ray radiation exhibited clear quasi-periodic modulations with a period of $\sim 15$ days (Pasham et al., 2024). The SMBH spin $a_{\bullet}\sim 0.25$ can be estimate from Fig.3. Moreover, the periodic X-ray emission shows two thermal components (a cool and warm component, and the temperature of thermal components also modulates on the same time scale, see Pasham et al., 2024), indicating that the line of sight aligns with the envelope’s inclination axis ($\theta_{\rm obs}\sim\theta_{\rm env}$). In this configuration, the observer can detect not only the thermal radiation from the inner thick disk within the envelope, but also the thermal emission produced by the cocoon shock breakout after the precessing jet is choked. If the cocoon shock breakout occurs on the other side (relative to the line of sight) of the ZEBRA surface, it would be obscured by the envelope, leading to periodic variations due to the jet precession [^2] Whether the cocoon shock breakout radiation at the envelope surface can be observed depends on the angle between the observer’s viewing direction ($\vec{r}_{\rm obs}$) and the normal of the surface element at the shock breakout region ($\vec{r}_{\rm co}$), only when $\vec{r}_{\rm obs}\cdot\vec{r}_{\rm co}>0$ (or the angle is less than $90^{\circ}$), the radiation can be observed.. In our model framework, the envelope of AT2020ocn has a large inclination angle, and the jet precession angle is smaller than the envelope inclination. We constrain the values of $\theta_{\rm obs}$ and $\theta_{\rm LT}$ for $\xi_{\rm duty}\sim 0.6$ in the panel (d) of Fig.5. Since only one constraint is applied, all points along the green solid line are potential solutions. Considering the significant precession expected from observations, we prefer a large $\theta_{\rm LT}$. Adopting the ZEBRA envelope parameters in the fiducial case, and assuming that the cocoon shock breakout is the Compton equilibrium between electrons and photons, one can calculate the cocoon breakout velocity $v_{\rm c,bo}\sim 0.1c$ with the observed temperature $T_{\rm c,bo}\sim 0.1~keV$ by the method in Zhu et al. (2021) and Nakar & Sari (2010). In the calculation, we use $\theta_{\rm j}\sim 5^{\circ}$, the breakout position $r_{\rm c,bo}\sim 0.9r_{\rm out}$ and density $\rho_{\rm c,bo}\sim 10^{-11}\rm g~cm^{-3}$ at this position. We can also rough estimate the value of the initial Lorentz factor ($\Gamma_{0}$) of the jet using the dynamical evolution equations for the jet and cocoon proposed by Huang et al. (1999) and Zhang et al. (2024) with the assumption of the jet choked position near the inner boundary of the envelope, and $\Gamma_{0}\sim 5$. Besides, similar modulations are not present in the optical bands (Pasham et al., 2024), which suggests that the optical radiation likely originates primarily from thermal emission at the envelope surface.
Figure 5: The estimation of $\theta_{\rm obs}$ and $\theta_{\rm LT}$ for Swift j1644+57 (panel a), AT 2022cmc (panel b), IGR J12580 (panel c) and AT 2020ocn (panel d) with the special $\xi_{\rm duty}$ or $\lambda$. The pink point in each panel represents the values of ($\theta_{\rm obs}$, $\theta_{\rm LT}$) that we adopt.
@@ 5 Conclusion
    The envelope density along its polar regions is significantly reduced, allowing the jets to escape freely from these funnels. For $\theta_{\rm LT}=\theta_{\rm env}$, observers with line of sight along these free zones can, therefore, detect a strong episodic jet, as the X-ray quasi-perodic dip shown in J1644. Away from these free zones, jets will be choked in most cases, and can only break out successfully at small precession angle. An off-axis jetted TDE, like J12580, is possible when looking along the choked zone.
@@
We are very grateful to Xiangyu Wang, and Rongfeng Shen for their helpful discussions. This work is supported by the National Natural Science Foundation of China under grants 12473012 and 12533005, and the National Key R&D Program of China (Nos. 2020YFC2201400, SQ2023YFC220007). W.H.Lei. acknowledges support by the science research grants from the China Manned Space Project with NO.CMS-CSST-2021-B11.